OpenStudy (anonymous):

FACTOR 2x^2 + x - 28 = 0 There should be 2 x's. I am stuck at 2(x^2+x-14)=0

OpenStudy (anonymous):

@haleyelizabeth2017 Lists steps

OpenStudy (haleyelizabeth2017):

Wait, it was originally \(2x^2+2x-28?\)

OpenStudy (haleyelizabeth2017):

What you have is the simplest it can go.

OpenStudy (anonymous):

Originally 2x^2+x-28=0

OpenStudy (haleyelizabeth2017):

Okay, then what you have so far is not correct.

OpenStudy (anonymous):

oh yea i made a mistake, i thought the GCF was 2 when it can't be

OpenStudy (anonymous):

what can be mult. to make -56 but added to make 1

OpenStudy (haleyelizabeth2017):

-7 and 4 multiply to get -28, okay? So then...if one side is (2x+/-_)(x+/-_) we can plug in those two numbers to try and see which works.

OpenStudy (anonymous):

im guessing that another way to do it Eli?

OpenStudy (haleyelizabeth2017):

So...let's try (2x-7)(x+4) first. We get \(2x^2+8x-7x-28\) which simplified, would work. So, (2x-7)(x+4) is the correct.

OpenStudy (haleyelizabeth2017):

correct answer* ^^

OpenStudy (anonymous):

so 2x-7=0 is x=7/2 or 3 1/2

OpenStudy (anonymous):

and x=-4

OpenStudy (haleyelizabeth2017):

Normally, it is left as an improper fraction, so x=7/2 is better. and x=-4 is correct as well.

OpenStudy (haleyelizabeth2017):

You could have also used the quadratic formula.

OpenStudy (haleyelizabeth2017):

It is a lot easier. Or, even complete the square works too

OpenStudy (haleyelizabeth2017):

\[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] \[ax^2+bx+c=0\]

OpenStudy (anonymous):

Ok, i got a question... the formula is ax^2+bx+c... i was told that you mult. a*c and the number that you get has to be mult. by x number and x number, but those two number must also be able to be added together to equal B... did that make sense

OpenStudy (anonymous):

Is there an exception to this??

OpenStudy (haleyelizabeth2017):

Huh?

OpenStudy (haleyelizabeth2017):

You don't have to do anything to the a b or c. just plug the coefficients into the quadratic formula and solve. Lets use your problem for example. We were given \[2x^2+x-28\] A would be 2, b would be 1, and c would be -28.

OpenStudy (haleyelizabeth2017):

\[\frac{ -(1) \pm \sqrt{(1)^2-4(2)(-28)} }{ 2(2) }\]\[\frac{ -1 \pm \sqrt{1+224} }{ 4 }\]and so on

OpenStudy (haleyelizabeth2017):

Did that make sense?

hero (hero):

@Beleaguer did you figure out What the two numbers are?

hero (hero):

In other words, can you solve m+n =1 mn =-56